# Opportunities and Challenges of Large Language Models for Low-Resource Languages in Humanities Research

### v3

@@ 4.1.1 Modeling of Low-Resource Languages in Cross-Cultural Contexts
Low-resource languages in cross-cultural contexts carry unique cultural values, often transmitting rich historical and social meanings through oral traditions, religious rituals, local ecologies, and artistic forms. Some oral traditions not only preserve the cultural semantics of ancestor worship but also shape intricate narrative rhythms. At the same time, the fusion of festive language with religious ritual language imbues regional languages with sacred elements in everyday usage. The integration of performing arts and architecture showcases the complex forms of language expression within cross-cultural settings[37]. Low-resource languages are often deeply influenced by these cultural factors, leading to complex and diverse modes of expression. Therefore, cross-cultural language modeling demands more than technical precision; it also requires integrating cultural understanding to develop models that reflect the cultural depth.
The corpora of low-resource languages, especially those carrying cultural histories such as oral traditions and folk narratives[129], are often insufficiently digitized or standardized, posing significant challenges for model adaptation. Some researchers have successfully incorporated these cultural differences into large models, yielding promising results[74]. In the processing of low-resource languages, studies have shown that domain-adaptive pretraining (DAPT) can significantly enhance the performance of language models on culturally specific corpora[96].
@@ 4.2.1 How Language Models Understand Polysemy Across Cultural Contexts
LLMs demonstrate a remarkable sensitivity and adaptability to cultural contexts when handling linguistic ambiguity. This is because the meaning of language is not only constrained by literal interpretation but is also deeply influenced by the cultural background and social norms in which it is used. Within the same language, specific words or expressions may exhibit different semantic layers due to cultural differences. By learning from vast, multilingual, and multicultural corpora, LLMs are capable of dynamically capturing contextual cues when analyzing text, enabling them to appropriately adapt to ambiguous words based on cultural context. This ability is particularly crucial in tasks such as language translation and cross-cultural dialogue generation, as it directly impacts the accuracy and cultural relevance of the generated language[98].
However, Polysemy - the existence of multiple but related meanings for a single form - has always been problematic for purely structural accounts of meaning[26]. LLMs face significant challenges when processing specific linguistic ambiguity - polysemy across different cultural contexts. The model’s performance largely depends on the breadth and diversity of its training data. If certain cultural corpora are underrepresented, or if the semantic characteristics of a particular culture are overrepresented, the model may provide overly simplified interpretations of ambiguous words, overlooking the subtle cultural differences underlying the language. The proposed computational framework for quantifying polysemy, as outlined in recent research, introduces a novel approach to tackling this challenge. By combining graph-based methods with syntactic structures, such as dependency parsing and Ollivier Ricci curvature, this framework enhances the model’s capacity to estimate polysemy scores and map linguistic ambiguity with precision[45]. Furthermore, linguistic ambiguity is often intertwined with metaphors, idioms, and other deep forms of expression, which poses even higher demands on the model. To enhance LLMs’ generalization ability in multicultural contexts, it is essential to incorporate balanced and multi-layered corpus resources, along with cultural context annotations and domain-specific knowledge to guide the model’s learning. These improvements can significantly enhance the model’s adaptability in multicultural environments, making it more accurate and comprehensive in language generation and understanding tasks.
@@ 4.2.2 Research on the Applicability of Language Models in Specific Cultural Scenarios
To enhance the applicability of LLMs in specific cultural environments, researchers can approach the issue from three key areas: data, architecture, and evaluation. First, integrating more localized and high-quality datasets into model training can help cover a wider range of cultural backgrounds and expressions[116]. Second, model architectures can include targeted modules, such as cultural context embeddings or bias correction mechanisms, to better accommodate diverse cultural needs. Lastly, by collaborating with local users and cultural experts, a cultural sensitivity evaluation system can be established to thoroughly test the model’s performance in specific cultural contexts. These efforts would contribute to promoting the fair and efficient use of LLMs in cross-cultural communication, education, content generation, and other fields.
@@ 4.3.1 Opportunities: Role of Low-Resource Language Models
Low-resource language models may have good potential in safeguarding endangered languages, traditions, and knowledge systems. These models empower the documentation and dissemination of oral traditions, folklore, and cultural narratives that might otherwise fade into obscurity due to a lack of balanced corpus resources. For example, the rich oral histories of indigenous communities, often passed down through generations, can be systematically transcribed and analyzed, ensuring their long-term preservation while simultaneously broadening their accessibility to a global audience. By bridging linguistic divides, language models[59] serve as conduits for cultural exchange, deepening cross-cultural intersectionality and fostering a sense of connection across diverse societies.
Beyond cultural preservation, low-resource language models offer a sophisticated method to unravel the complexities inherent in cross-cultural communication[75]. These models are particularly adept at addressing linguistic differences, such as dialectal variations, where they can act as multilingual experts, bridging and connecting diverse linguistic systems. With the potential to facilitate seamless translation and communication across languages and dialects without delays or loss of authenticity, language models are redefining the scope of cross-cultural interactions. By fine-tuning these models on culturally specific corpora, they can capture nuanced meanings with remarkable accuracy within the specific area, fostering a deeper understanding of the source culture while ensuring that translations and interpretations faithfully reflect original intention and the cultural depth.
@@ 4.3.2 Challenges: Misunderstandings and Bias within Low-Resource Language Models
While the potential of low-resource language models in advancing cross-cultural communication and preserving endangered traditions is undeniably promising, several significant obstacles stand in the way. A major challenge is the risk of misunderstandings and biases arising from insufficient cultural and contextual understanding. Currently, many language models are developed using training datasets that fail to fully encompass the linguistic diversity and cultural specificity of the communities they represents. As a result, the nuanced meanings of idiomatic expressions, metaphors, and culturally embedded terms are often misunderstood within cross-cultural context, leading to oversimplified translations or inaccuracies that fail to honor the richness of the original text[79].
Another pressing challenge lies in the inadvertent projection of dominant cultural perspectives onto minority languages or traditions. Many LLMs are predominantly trained on datasets drawn from high-resource languages and cultures, leading to a biased interpretative framework. This bias can dilute or distort the unique cultural and linguistic attributes of minority traditions, reducing their authenticity and misrepresenting their core meanings. Moreover, the scarcity of comprehensive resources for many low-resource languages exacerbates this issue. Without access to curated datasets and the necessary external contextual resources, LLMs struggle to accurately interpret and convey these elements. This challenge is particularly acute for oral traditions and languages with limited documentation, where much of the cultural essence remains undocumented or inaccessible[68]. Taking these and other unknown limitations into consideration, there is a long way to achieve the effective application for low-resource language models in cross-cultural context.
@@ 5.1 Automated Translation and Analysis of Low-Resource Literary Works
Literary works play a crucial role in the language system. Literary works are not only the simple transmission and preservation of information, but they also carry the author’s emotions, thoughts, and even contain the cultural background of a region or even a country[103]. Therefore, automated translation and analysis of literary works is an important way to understand and inherit a language[65]. The translation of literary works is not just about simple word conversion, but about conveying the same emotions and meanings as the original work in a new language environment. Similarly, the analysis of literary works also needs to uncover the deep hidden meanings of the works, rather than simply retelling and summarizing them. For low-resource languages such as Croatian and most other less used languages[33], it is crucial to use LLMs to translate and analyze literary works for understanding, protecting, and even cultural heritage.
However, using LLMs to translate and analyze literary works in low-resource languages is a relatively difficult process. This is because for low-resource languages, there is usually a lack of large-scale parallel corpora or high-quality datasets[78], which are the basis for LLMs to understand low-resource languages, making it difficult to process literary works. Secondly, literary works in low-resource languages often contain unique cultural backgrounds and local characteristics. Without sufficient cultural background knowledge, it is difficult for LLMs to accurately translate these unique cultural elements. To address these challenges, researchers have proposed various methods, including building corpora of low resource languages, cross language transfer learning[56], and using professionals to proofread and revise the results provided by large language models.
@@ 5.2 Identification of Specific Vocabulary and Concepts in Religious Texts
Religious texts are usually written in low-resource languages and are important repository of theological, philosophical, and cultural knowledge. They are renowned for their unique terminology, symbolic structures, and contextual meanings[2], and can be better studied and analyzed using LLMs.
@@
Lai’s research[71] offers an interdisciplinary approach to linguistics, combining qualitative theory (Dialogic Syntax, Construction Grammar) with quantitative computational analysis focused on corpus linguistics, statistical modeling, and information theory. This ”usage-based” perspective views language as a dynamic system emerging from human interaction, exemplified by the analysis of texts like the Diamond Sutra. He argues that highly structured, non-spontaneous texts, like orally transmitted religious texts, provide a valuable, high-signal source for understanding linguistic patterns, suggesting a novel approach to LLM training by curating stylized, ritualized texts for enhanced structural and relational richness.
@@
Religious texts generally have specific language and structural attributes. Many low-resource religious texts use vocabulary with profound theological significance, often lacking direct equivalents in modern language, which are referred to as sacred terminology. For example, the Sanskrit term ”Nirvana” [108]represents a complex concept that differs greatly between Hinduism and Buddhism, including liberation, cessation, and transcendence; In Christian texts, Greek words like ”agape” and ”logos” are filled with theological and philosophical meanings. In addition, some religious texts may use repetitive phrases for ritual, memory, or meditation purposes, and these repetitions need to be explained and preserved during the analysis and translation process[36]. Secondly, many religious texts heavily rely on symbolism and fables[60]. For example, the recurring theme of light in sacred texts such as Vedic hymns or the Bible conveys a range of meanings from divine existence to enlightenment. Finally, religious texts are deeply rooted in their social and historical context. Words, phrases, and even entire paragraphs can reflect specific historical events, cultural norms, or local belief systems, and require contextual understanding to be accurately interpreted.
@@
The LLMs provide revolutionary possibilities[5] for the study of religious texts. In cross religious contexts, LLMs can compare and study traditional texts of different religions, identify terminology and doctrines. For example, LLMs can compare the concept of ”sympathy” expressed in Buddhist Pali scriptures, Christian New Testament texts, and Islamic hadith, providing insights into both general and specific traditional interpretations. LLMs can help explain the subtle differences of the same term in different religious backgrounds. For example, the Sanskrit term ”dharma” [15] in Hinduism may refer to justice or responsibility, while in Buddhism, it means the teachings of the Buddha or universal truth. LLMs trained extensively in context can help eliminate ambiguity in these terms during analysis or translation processes. In the cross-religious context, the same symbol may have different meanings in different traditions. Religious and philosophical texts are not simply repositories of information; they are highly structured performances of knowledge and authority, featuring unique rhetorical and structural patterns that are not present in everyday communication. A simple factual summary of such texts fails to capture their underlying rhetorical and dialogic function. For example, the lion is a symbol of strength and divinity in the Judeo Christian tradition, while in the Buddhist tradition it represents wisdom and courage. LLMs can compare these symbols in the corpus and enrich the interpretation by placing them in their respective traditions.
@@ 5.3.1 Opportunities: Innovative Applications of Language Models
To revive low-resource literary works and ensure their resonance with modern audiences, particularly younger generations, the presentation and dissemination of these literary works must align with contemporary tastes and cultural dynamics. Language models can facilitate this by economically producing accessible, high-quality, and readable content tailored to the preferences of target audiences. Leveraging popular platforms and modern formats, such as interactive media or engaging digital narratives, can reintroduce overlooked literary treasures in a way that is both appealing and relevant. This approach not only preserves the vibrancy and influence of these cultural artifacts but also fosters a deeper understanding of global heritage while enabling meaningful dialogue on religious and cultural complexities. Specifically, it helps younger readers engage more effectively with religious texts, improving theological comprehension[72] and contributing to the movement of intercultural peace.
@@
Beyond accessibility and revitalization, LLMs can also be a powerful tool for academic and theological inquiry. In the realm of religious studies, language models offer a unique opportunity to address doctrinal conflicts and enrich theological discourse in several areas. By analyzing religious texts, these models can uncover common principles and provide clarity in areas where doctrinal interpretations have historically diverged. From a technological perspective, the systematically machine-learnable structure inherent in the Buddha-Subhti dialogue, as developed by Ryan Lai, presents several contributions to the fields of AI[70] and low-resource language processing. These contributions include the comprehensive modeling of epistemic status, the operationalization of rhetorical routines, and the recognition of the inherent value of stylized corpora for religious text, which furnish high-quality training data due to their structured textual nature. Through a profound understanding of linguistic routines and their associated social functions, this research endeavors to facilitate the development of more nuanced and communicatively proficient LLMs, particularly within intricate and specialized domains.
@@
This article, alongside that of scholars like Carolyn Medine, underscores a crucial point: the successful application of LLMs in religious studies requires a comprehensive understanding of both the linguistic structure of sacred texts and their broader social and cultural contexts. For example, a purely textual analysis of Buddhism is insufficient without considering how entrenched racial and cultural dynamics shape its interpretations, as examined in [141]. Likewise, Carolyn Medine investigates the interdisciplinary culture of race, discrimination, and religion[97] in her tribute article to bell hooks, a black feminist and black buddhist, it highlights the need for LLMs to incorporate the nuances of race, gender, and social context in their analysis beyond a purely textual approach. Therefore, these models are capable of adapting and refining doctrinal interpretations to align with evolving perspectives, while remaining anchored in traditional frameworks. A notable application is the preservation and dissemination of religious traditions tied to specific cultural contexts, such as African-American theology[62]. With thoughtful development and ethical oversight, these models have the capacity to establish balanced frameworks that address doctrinal ambiguities while expanding the accessibility of religious knowledge across diverse cultural boundaries.

### v4

@@ 4.1.1 Modeling of Low-Resource Languages in Cross-Cultural Contexts
Low-resource languages in cross-cultural contexts carry unique cultural values, often transmitting rich historical and social meanings through oral traditions, religious rituals, local ecologies, and artistic forms. Some oral traditions not only preserve the cultural semantics of ancestor worship but also shape intricate narrative rhythms. At the same time, the fusion of festive language with religious ritual language imbues regional languages with sacred elements in everyday usage. The integration of performing arts and architecture showcases the complex forms of language expression within cross-cultural settings[erasmo2020theatre]. Low-resource languages are often deeply influenced by these cultural factors, leading to complex and diverse modes of expression. Therefore, cross-cultural language modeling demands more than technical precision; it also requires integrating cultural understanding to develop models that reflect the cultural depth.
The corpora of low-resource languages, especially those carrying cultural histories such as oral traditions and folk narratives[tehrani2023cultural], are often insufficiently digitized or standardized, posing significant challenges for model adaptation. Some researchers have successfully incorporated these cultural differences into large models, yielding promising results[li2024culturellm]. In the processing of low-resource languages, studies have shown that domain-adaptive pretraining (DAPT) can significantly enhance the performance of language models on culturally specific corpora[meaney2024evaluating].
@@ 4.2.1 How Language Models Understand Polysemy Across Cultural Contexts
LLMs demonstrate a remarkable sensitivity and adaptability to cultural contexts when handling linguistic ambiguity. This is because the meaning of language is not only constrained by literal interpretation but is also deeply influenced by the cultural background and social norms in which it is used. Within the same language, specific words or expressions may exhibit different semantic layers due to cultural differences. By learning from vast, multilingual, and multicultural corpora, LLMs are capable of dynamically capturing contextual cues when analyzing text, enabling them to appropriately adapt to ambiguous words based on cultural context. This ability is particularly crucial in tasks such as language translation and cross-cultural dialogue generation, as it directly impacts the accuracy and cultural relevance of the generated language[kibria-etal-2024-functional].
However, Polysemy - the existence of multiple but related meanings for a single form - has always been problematic for purely structural accounts of meaning[DEANE1988325]. LLMs face significant challenges when processing specific linguistic ambiguity - polysemy across different cultural contexts. The model’s performance largely depends on the breadth and diversity of its training data. If certain cultural corpora are underrepresented, or if the semantic characteristics of a particular culture are overrepresented, the model may provide overly simplified interpretations of ambiguous words, overlooking the subtle cultural differences underlying the language. The proposed computational framework for quantifying polysemy, as outlined in recent research, introduces a novel approach to tackling this challenge. By combining graph-based methods with syntactic structures, such as dependency parsing and Ollivier Ricci curvature, this framework enhances the model’s capacity to estimate polysemy scores and map linguistic ambiguity with precision[Goel2023BeyondTS]. Furthermore, linguistic ambiguity is often intertwined with metaphors, idioms, and other deep forms of expression, which poses even higher demands on the model. To enhance LLMs’ generalization ability in multicultural contexts, it is essential to incorporate balanced and multi-layered corpus resources, along with cultural context annotations and domain-specific knowledge to guide the model’s learning. These improvements can significantly enhance the model’s adaptability in multicultural environments, making it more accurate and comprehensive in language generation and understanding tasks.
@@ 4.2.2 Research on the Applicability of Language Models in Specific Cultural Scenarios
To enhance the applicability of LLMs in specific cultural environments, researchers can approach the issue from three key areas: data, architecture, and evaluation. First, integrating more localized and high-quality datasets into model training can help cover a wider range of cultural backgrounds and expressions[rao2024normad]. Second, model architectures can include targeted modules, such as cultural context embeddings or bias correction mechanisms, to better accommodate diverse cultural needs. Lastly, by collaborating with local users and cultural experts, a cultural sensitivity evaluation system can be established to thoroughly test the model’s performance in specific cultural contexts. These efforts would contribute to promoting the fair and efficient use of LLMs in cross-cultural communication, education, content generation, and other fields.
@@ 4.3.1 Opportunities: Role of Low-Resource Language Models
Low-resource language models may have good potential in safeguarding endangered languages, traditions, and knowledge systems. These models empower the documentation and dissemination of oral traditions, folklore, and cultural narratives that might otherwise fade into obscurity due to a lack of balanced corpus resources. For example, the rich oral histories of indigenous communities, often passed down through generations, can be systematically transcribed and analyzed, ensuring their long-term preservation while simultaneously broadening their accessibility to a global audience. By bridging linguistic divides, language models[Hutson2024] serve as conduits for cultural exchange, deepening cross-cultural intersectionality and fostering a sense of connection across diverse societies.
Beyond cultural preservation, low-resource language models offer a sophisticated method to unravel the complexities inherent in cross-cultural communication[li2024cultureparkboostingcrossculturalunderstanding]. These models are particularly adept at addressing linguistic differences, such as dialectal variations, where they can act as multilingual experts, bridging and connecting diverse linguistic systems. With the potential to facilitate seamless translation and communication across languages and dialects without delays or loss of authenticity, language models are redefining the scope of cross-cultural interactions. By fine-tuning these models on culturally specific corpora, they can capture nuanced meanings with remarkable accuracy within the specific area, fostering a deeper understanding of the source culture while ensuring that translations and interpretations faithfully reflect original intention and the cultural depth.
@@ 4.3.2 Challenges: Misunderstandings and Bias within Low-Resource Language Models
While the potential of low-resource language models in advancing cross-cultural communication and preserving endangered traditions is undeniably promising, several significant obstacles stand in the way. A major challenge is the risk of misunderstandings and biases arising from insufficient cultural and contextual understanding. Currently, many language models are developed using training datasets that fail to fully encompass the linguistic diversity and cultural specificity of the communities they represents. As a result, the nuanced meanings of idiomatic expressions, metaphors, and culturally embedded terms are often misunderstood within cross-cultural context, leading to oversimplified translations or inaccuracies that fail to honor the richness of the original text[liu2023wereafraidlanguagemodels].
Another pressing challenge lies in the inadvertent projection of dominant cultural perspectives onto minority languages or traditions. Many LLMs are predominantly trained on datasets drawn from high-resource languages and cultures, leading to a biased interpretative framework. This bias can dilute or distort the unique cultural and linguistic attributes of minority traditions, reducing their authenticity and misrepresenting their core meanings. Moreover, the scarcity of comprehensive resources for many low-resource languages exacerbates this issue. Without access to curated datasets and the necessary external contextual resources, LLMs struggle to accurately interpret and convey these elements. This challenge is particularly acute for oral traditions and languages with limited documentation, where much of the cultural essence remains undocumented or inaccessible[kirk2023personalisationboundsrisktaxonomy]. Taking these and other unknown limitations into consideration, there is a long way to achieve the effective application for low-resource language models in cross-cultural context.
@@ 5.1 Automated Translation and Analysis of Low-Resource Literary Works
Literary works play a crucial role in the language system. Literary works are not only the simple transmission and preservation of information, but they also carry the author’s emotions, thoughts, and even contain the cultural background of a region or even a country[olsen1982meaning]. Therefore, automated translation and analysis of literary works is an important way to understand and inherit a language[karabayeva2024evaluating]. The translation of literary works is not just about simple word conversion, but about conveying the same emotions and meanings as the original work in a new language environment. Similarly, the analysis of literary works also needs to uncover the deep hidden meanings of the works, rather than simply retelling and summarizing them. For low-resource languages such as Croatian and most other less used languages[dundjer2020automatic], it is crucial to use LLMs to translate and analyze literary works for understanding, protecting, and even cultural heritage.
However, using LLMs to translate and analyze literary works in low-resource languages is a relatively difficult process. This is because for low-resource languages, there is usually a lack of large-scale parallel corpora or high-quality datasets[lin2020towards], which are the basis for LLMs to understand low-resource languages, making it difficult to process literary works. Secondly, literary works in low-resource languages often contain unique cultural backgrounds and local characteristics. Without sufficient cultural background knowledge, it is difficult for LLMs to accurately translate these unique cultural elements. To address these challenges, researchers have proposed various methods, including building corpora of low resource languages, cross language transfer learning[huang2020cross], and using professionals to proofread and revise the results provided by large language models.
@@ 5.2 Identification of Specific Vocabulary and Concepts in Religious Texts
Religious texts are usually written in low-resource languages and are important repository of theological, philosophical, and cultural knowledge. They are renowned for their unique terminology, symbolic structures, and contextual meanings[agliz2015translation], and can be better studied and analyzed using LLMs.
@@
Lai’s research[lai-etal-2023-turn] offers an interdisciplinary approach to linguistics, combining qualitative theory (Dialogic Syntax, Construction Grammar) with quantitative computational analysis focused on corpus linguistics, statistical modeling, and information theory. This ”usage-based” perspective views language as a dynamic system emerging from human interaction, exemplified by the analysis of texts like the Diamond Sutra. He argues that highly structured, non-spontaneous texts, like orally transmitted religious texts, provide a valuable, high-signal source for understanding linguistic patterns, suggesting a novel approach to LLM training by curating stylized, ritualized texts for enhanced structural and relational richness.
@@
Religious texts generally have specific language and structural attributes. Many low-resource religious texts use vocabulary with profound theological significance, often lacking direct equivalents in modern language, which are referred to as sacred terminology. For example, the Sanskrit term ”Nirvana” [pasadika2007nirvana]represents a complex concept that differs greatly between Hinduism and Buddhism, including liberation, cessation, and transcendence; In Christian texts, Greek words like ”agape” and ”logos” are filled with theological and philosophical meanings. In addition, some religious texts may use repetitive phrases for ritual, memory, or meditation purposes, and these repetitions need to be explained and preserved during the analysis and translation process[elewa2014features]. Secondly, many religious texts heavily rely on symbolism and fables[jakel2002hypotheses]. For example, the recurring theme of light in sacred texts such as Vedic hymns or the Bible conveys a range of meanings from divine existence to enlightenment. Finally, religious texts are deeply rooted in their social and historical context. Words, phrases, and even entire paragraphs can reflect specific historical events, cultural norms, or local belief systems, and require contextual understanding to be accurately interpreted.
@@
The LLMs provide revolutionary possibilities[Sacred_Texts_Digital_Age] for the study of religious texts. In cross religious contexts, LLMs can compare and study traditional texts of different religions, identify terminology and doctrines. For example, LLMs can compare the concept of ”sympathy” expressed in Buddhist Pali scriptures, Christian New Testament texts, and Islamic hadith, providing insights into both general and specific traditional interpretations. LLMs can help explain the subtle differences of the same term in different religious backgrounds. For example, the Sanskrit term ”dharma” [brockington2004concept] in Hinduism may refer to justice or responsibility, while in Buddhism, it means the teachings of the Buddha or universal truth. LLMs trained extensively in context can help eliminate ambiguity in these terms during analysis or translation processes. In the cross-religious context, the same symbol may have different meanings in different traditions. Religious and philosophical texts are not simply repositories of information; they are highly structured performances of knowledge and authority, featuring unique rhetorical and structural patterns that are not present in everyday communication. A simple factual summary of such texts fails to capture their underlying rhetorical and dialogic function. For example, the lion is a symbol of strength and divinity in the Judeo Christian tradition, while in the Buddhist tradition it represents wisdom and courage. LLMs can compare these symbols in the corpus and enrich the interpretation by placing them in their respective traditions.
@@ 5.3.1 Opportunities: Innovative Applications of Language Models
To revive low-resource literary works and ensure their resonance with modern audiences, particularly younger generations, the presentation and dissemination of these literary works must align with contemporary tastes and cultural dynamics. Language models can facilitate this by economically producing accessible, high-quality, and readable content tailored to the preferences of target audiences. Leveraging popular platforms and modern formats, such as interactive media or engaging digital narratives, can reintroduce overlooked literary treasures in a way that is both appealing and relevant. This approach not only preserves the vibrancy and influence of these cultural artifacts but also fosters a deeper understanding of global heritage while enabling meaningful dialogue on religious and cultural complexities. Specifically, it helps younger readers engage more effectively with religious texts, improving theological comprehension[leblanc2012ancient] and contributing to the movement of intercultural peace.
@@
Beyond accessibility and revitalization, LLMs can also be a powerful tool for academic and theological inquiry. In the realm of religious studies, language models offer a unique opportunity to address doctrinal conflicts and enrich theological discourse in several areas. By analyzing religious texts, these models can uncover common principles and provide clarity in areas where doctrinal interpretations have historically diverged. From a technological perspective, the systematically machine-learnable structure inherent in the Buddha-Subhti dialogue, as developed by Ryan Lai, presents several contributions to the fields of AI[lai-2024-tupleised] and low-resource language processing. These contributions include the comprehensive modeling of epistemic status, the operationalization of rhetorical routines, and the recognition of the inherent value of stylized corpora for religious text, which furnish high-quality training data due to their structured textual nature. Through a profound understanding of linguistic routines and their associated social functions, this research endeavors to facilitate the development of more nuanced and communicatively proficient LLMs, particularly within intricate and specialized domains.
@@
This article, alongside that of scholars like Carolyn Medine, underscores a crucial point: the successful application of LLMs in religious studies requires a comprehensive understanding of both the linguistic structure of sacred texts and their broader social and cultural contexts. For example, a purely textual analysis of Buddhism is insufficient without considering how entrenched racial and cultural dynamics shape its interpretations, as examined in [yancy2019buddhism]. Likewise, Carolyn Medine investigates the interdisciplinary culture of race, discrimination, and religion[Medine_2022] in her tribute article to bell hooks, a black feminist and black buddhist, it highlights the need for LLMs to incorporate the nuances of race, gender, and social context in their analysis beyond a purely textual approach. Therefore, these models are capable of adapting and refining doctrinal interpretations to align with evolving perspectives, while remaining anchored in traditional frameworks. A notable application is the preservation and dissemination of religious traditions tied to specific cultural contexts, such as African-American theology[jones_medine_2023]. With thoughtful development and ethical oversight, these models have the capacity to establish balanced frameworks that address doctrinal ambiguities while expanding the accessibility of religious knowledge across diverse cultural boundaries.